In a game of chance a woman is paid $ 15 if she draws a card of hearts, and $ 5 if she draws a card of diamonds from an ordinary deck of 52 cards (an ordinary deck has 13 cards of hearts, 13 of diamonds, 13 of clovers and 13 spades). If you draw any other card, you lose $ 13. If she plays, what would be the expected profit of the woman?

a) -$0.75

b) -$1.5

c) -$1.25

d) -$0.25
Answer #1

Problem in this then comment below.. i will help you..

.

please thumbs up for this solution... Thanks..

.

expected value = 15/4 +5/4 - 13/2 = -3/2

.

answer = option b ... - $ 1.5